OpenStudy (rissa):

The mosquito population within a defined area is estimated at 45 million. Due to weather conditions, it is projected that this population will increase by 40%. Find the increase and the projected population.

OpenStudy (a_clan):

increase=40% of 45 m =(40/100)45m =18,000,000 = 18 m projected = 45m + 18m = 63m